# AL-Bench: A Benchmark for Automatic Logging

### v2

@@ 4. EXPERIMENTS
  • RQ1: How well are logging tools at predicting log statements in static evaluation?
  • RQ2: How well can the predicted log statements be compiled successfully?
  • RQ3: How well do the generated log statements print logs in dynamic evaluation?
Specifically, RQ1 examines how well logging tools predict log statements using the AL-Bench static evaluation dataset and corresponding static evaluation method. This research question aims to evaluate the accuracy and effectiveness of the tools in generating log statements that closely resemble the ground truth. RQ2 focuses on assessing the compilability of the predicted log statements, determining whether these predictions can be seamlessly integrated into the code without causing compilation errors. Finally, RQ3 evaluates the runtime logs produced by the generated log statements. We examine not only the generation of inappropriate logs but also the similarity between the predicted logs and the expected logs in cases where logs are correctly generated. We begin by introducing the automatic logging tools selected for evaluation. We will provide a detailed analysis of these tools, focusing on their performance across the three research questions. Each tool is assessed for its ability to predict accurate log statements, compile them successfully, and generate meaningful logs. This comprehensive evaluation allows us to draw insights into the strengths and limitations of current approaches to automatic logging.
@@ 4.1. Automatic Logging Tools
Automatic logging is a hot topic, leading to the development of many tools for determining specific parts of log statements and end-to-end automatic logging tools in recent years. In this paper, we focus on end-to-end logging tools. We reached out to the authors of popular end-to-end automatic logging tools for assistance in rebuilding these tools. Because SCLogger (Li et al., 2024) is still under construction, we ultimately selected four methods for evaluation: LANCE (Mastropaolo et al., 2022), LEONID (Mastropaolo et al., 2023), FastLog (Xie et al., 2023), UniLog (Xu et al., 2024). We detailed the method in the following.
LANCE: $LANCE$ (Mastropaolo et al., 2022) is the first model designed to generate and insert complete log statements in code. It takes a method requiring a log statement and outputs a meaningful log message with an appropriate logging level in the correct position. Built on the Text-To-Text Transfer Transformer model, $LANCE$ is trained specifically for injecting proper logging statements.
@@
UniLog: $UniLog$ (Xu et al., 2024) is the first attempt to adapt Warm-up and In-context-learning strategy to enhance the model’s ability to generate log statements. Due to limitations in assessing the original UniLog, we reproduced it using two backbone models: CodeLlama-7B (Rozière et al., 2023) and DeepSeek-V3 (DeepSeek-AI et al., 2024). We applied the warmup process exclusively to the CodeLlama backbone model, while employing the ICL strategy to construct prompts for both models. The data are sourced from LANCE (Mastropaolo et al., 2022) to warm up and generate ICL content. The effectiveness of In-Context Learning often depends on whether the examples are in-distribution or out-of-distribution relative to the evaluation data. Since LANCE’s data distribution differs from our evaluation data, this may affect UniLog’s performance. We will use $UniLog_{{cl}}$ to represent the version based on CodeLlama-7B, $UniLog_{{ds}}$ to represent the version based on DeepSeek-V3.
@@
FastLog: FastLog (Xie et al., 2023) defines the logging task in two steps: finding the position and generating and inserting a complete log statement into the source code. This approach avoids rewriting the source code, a key limitation of $LANCE$. They utilized PLBART (Ahmad et al., 2021) as the base model to fine-tune two separate models: one for predicting insertion position, the other for generating log statements. With the heuristic rule, log statements only appear after certain special characters, FastLog enhances efficiency while maintaining accuracy in generating log statements.
Table 2. Static evaluation on the complete logging task was conducted. The values in the Original lines indicate the reported results from the respective methods’ own evaluations. The best performance evaluated by AL-Bench across different metrics is highlighted.
Method PA LA MA DVA STB (BLEU)
Now Original Now Original Now Original
FastLog 57.54 $58.84$ 62.72 $59.75$ 6.90 $4.52$ 17.66 20.20
$UniLog_{{ds}}$ $36.31$ $76.90$ $36.31$ $72.30$ $5.19$ $22.40$ 15.73 11.60
$UniLog_{{cl}}$ $23.29$ $76.90$ $23.29$ $72.30$ $2.31$ $22.40$ 16.08 8.43
LANCE $34.67$ $65.40$ $34.67$ $66.24$ $2.99$ $16.90$ 14.42 6.52
$LEONID_{S}$ $15.22$ $76.45$ $15.22$ $73.53$ $1.38$ $31.55$ 5.35 2.03
4.2. RQ1: How well are logging tools at predicting log statements in static evaluation?
@@
We evaluated four automatic logging tools using the AL-Bench static evaluation dataset and corresponding static evaluation methods, ensuring a fair comparison of their performance.
Following previous studies (Mastropaolo et al., 2022; Xu et al., 2024; Xie et al., 2023; Mastropaolo et al., 2023), we first processed our dataset to create evaluation pairs $<M_{s},M_{t}>$ with $M_{s}$ representing the input provided to the model (i.e., $M_{s}$ with one removed log statement) and $M_{t}$ being the expected output (i.e., $M_{t}$ is the removed log statement). After processing, we got 42,224 instances. The evaluation results are presented in Table 2. The metrics for Position Accuracy (PA), Level Accuracy (LA), and Message Accuracy (MA) are adapted from the evaluation frameworks used by the selected logging tools. Therefore, we label the values of these metrics as reported in their respective evaluation studies. The results demonstrate significant performance differences among the four logging tools.
FastLog consistently outperforms the other tools across all metrics, but it still faces challenges in specific areas. For instance, it scores only 17.66 in Dynamic Variable Accuracy (DVA) and 20.44 in Static Text BLEU (STB), suggesting that even the best tool struggles to generate meaningful natural language descriptions and accurately select dynamic variables for logging. Despite these shortcomings, FastLog shows only minor deviations from previous studies, maintaining stable performance across all metrics. In contrast, the other three tools, particularly LEONID, experience a significant decline in accuracy across all evaluated metrics.
As illustrated in Figure 7, the performance of each tool varies considerably across different projects. This fluctuation suggests that the reliability and stability of logging tools remain inconsistent when applied to diverse scenarios. The performance fluctuations of logging tools across different projects are due to variations in project complexity and the diverse logging requirements. Additionally, limited training data diversity may hinder the tools’ ability to perform consistently across unfamiliar scenarios. These factors collectively contribute to the observed instability in logging tool performance across different projects. As demonstrated in our results, the variability in tool performance across different projects underscores the necessity of diverse evaluation data to capture the full spectrum of a tool’s capabilities and limitations.
To investigate the causes of performance differences compared to previous studies, we observed that earlier evaluation datasets were filtered to exclude instances longer than 512 tokens. To assess the impact of this filtering, we divided our dataset into two groups: instances with shorter than 512 tokens and those with longer than 512 tokens. This division enables us to determine whether instance length contributes to the performance discrepancies observed among the logging tools, providing further insight into how different data characteristics influence tool effectiveness.
Table 3. The performance of tools when facing different length input data. Long means data longer than 512 tokens, Short means data shorter than 512 tokens, and $\Delta$ means the difference in logging tools performance when facing the Short data and the Long data.
Method Dataset PA LA MA DVA STB
FastLog Long 43.36 64.92 7.02 11.98 20.97
Short 61.79 62.33 6.88 19.36 19.98
$\Delta$ $\downarrow 18.43$ $\uparrow 2.59$ $\uparrow 0.14$ $\downarrow 7.38$ $\uparrow 0.99$
$UniLog_{d}{}_{s}$ Long 22.28 66.66 8.82 15.56 17.05
Short 40.51 59.52 4.11 15.78 9.97
$\Delta$ $\downarrow 18.23$ $\uparrow 7.14$ $\uparrow 4.11$ $\downarrow 0.22$ $\uparrow 7.08$
$UniLog_{c}{}_{l}$ Long 8.70 51.62 2.84 12.91 8.20
Short 27.67 51.78 2.62 17.02 8.80
$\Delta$ $\downarrow 18.97$ $\downarrow 0.16$ $\uparrow 0.22$ $\downarrow 4.12$ $\downarrow 0.60$
LANCE Long 0 0 0 0 0
Short 45.05 47.61 3.89 18.74 8.12
$\Delta$
$LEONID_{S}$ Long 0 0 0 0 0
Short 19.78 21.46 1.79 6.95 2.64
$\Delta$
After dividing the dataset, we obtained two groups: a longer dataset containing 9,732 instances and a shorter dataset with 32,492 instances. The evaluation results in Table 3 show a clear difference in tool performance based on the length of the instances. Notably, LANCE and LEONID struggled with instances longer than 512 tokens, failing to generate syntax-correct code and, in some cases, producing incomplete code. This explains why their scores for these cases are reported as zero, highlighting the input length limitations of both tools and their inability to handle longer, more complex instances effectively. $UniLog_{cl}$ and FastLog show a considerable drop in PA and DVA when handling longer data, indicating that they struggle to predict log positions and select dynamic variables in more complex instances. $UniLog_{ds}$also shows a dramatic drop in position accuracy, but only a slight decrease in the ability to decide dynamic variables for recording. However, with more sufficient information and a more powerful backbone model, it demonstrates significantly better performance in other metrics. This indicates that with a more powerful backbone, UniLog can better understand the context with longer inputs. Although $UniLog_{d}s$ is equipped with a more powerful backbone, it still struggles to choose the appropriate position for the log statement. As widely demonstrated that the LLMs are not good at counting numbers (Ahn et al., 2024), we might change the output format by directly adding log statements into the source code rather than generating the exact line number. Furthermore, by simply analyzing the control flow graph of the code, we might be able to exclude positions where logging is not feasible to leave fewer choices.
Answer to RQ1. FastLog performs best in AL-Bench ’s static evaluation but still faces challenges in generating meaningful descriptions and selecting key dynamic variables. All tools struggle to maintain stable performance under varying logging requirements and complex data, especially in determining log positions. Since LLMs struggle with counting, it might be better to avoid outputting exact line numbers and instead add tags in the source code. Additionally, analyzing the control graph to exclude impossible positions might be useful to enhance log position determination.
@@
Failed Reason Failed Number
@@
4.3. RQ2: How well can the predicted log statements be compiled successfully?
To evaluate the tools’ ability to generate compilable log statements, we replace existing logs with predicted ones in the dynamic evaluation dataset and recompile the project to check for successful compilation. As shown in Table 5, the best-performing tool, FastLog, achieves a 79.9% Compilation Success Rate, followed by $UniLog_{c}l$ at 70.3%, $UniLog_{d}s$ at 60.2%, LANCE at 49.4%, $LEONID_{M}$ and $LEONID_{S}$ at 25.0% and 16.4%, respectively. The results reveal a key limitation of LANCE and LEONID that they regenerate the entire code snippet, which increases the risk of unintended code changes and can potentially lead to compilation errors. In contrast, FastLog and UniLog focus solely on generating the new log statement, minimizing the risk of errors by limiting codebase modifications. Although FastLog is one of the best tools available, it still leads to significant instances of compile failure. When log statements fail to compile, they can prevent the system from functioning as intended and complicate debugging or error tracking. To ensure the reliability of automatic logging tools in real-world environments, it’s crucial that the log statements they generate integrate smoothly into the existing codebase.
Given the importance of reliable compilation, we conducted a manual review of the compilation failures to identify the specific causes. This analysis will provide further insight into the key weaknesses of these tools and potential areas for improvement. We randomly selected 100 failed instances from the best tool, FastLog, and the first two authors cross-checked the causes. The analysis results are presented in Table 5. The most common failure, occurring in 56 instances, is due to Using the wrong Logging Name (i.e., Logger, LOG), indicating that incorrect or non-existent log functions are being invoked. Using Undefined Methods accounts for 21 failures, followed by Using Undefined Variables with 15 failures. Less frequent issues include Incompatible Types (3), and Unreachable Statements (1). Others mean generating the wrong syntax code, which we will not analyze.
The majority of failures (totaling 92 instances) involve undefined references to methods (21 instances), variables (15 instances), or logging names (56 instances). These undefined reference failures are primarily due to the limited context provided by existing function-level logging tools, lacking critical details on valid variables, methods, libraries, and packages relevant to the target function. Current tools are designed for function-level input, highlighting the need for logging tools to integrate better context awareness and validation checks to ensure compatibility with the existing codebase. The less frequent errors, such as incompatible types and unreachable statements, also indicate challenges in generating log statements that integrate well with the surrounding code logic. The issue of Unreachable Statements is notable because it points to a fundamental weakness in the current logging tools—specifically, their lack of understanding of the code’s control flow.
@@ 4.3. RQ2: How well can the predicted log statements be compiled successfully?
4.4. RQ3: How well do the generated log statements print logs in dynamic evaluation?
To answer RQ3, we used the dynamic evaluation dataset including code snippets paired with corresponding unit tests to generate logs. For each instance, we replaced the existing log statements in the source code with the predicted log statements from the automatic logging tools. The unit tests were then executed to observe the logs generated by the modified code.
It is important to highlight that, aside from $LEONID_{M}$, all other tools operate under the strong assumption that the given code snippet requires exactly one log statement. To highlight this inappropriate assumption, we only allow the tools one chance to predict the log statement, even in cases where multiple log statements might be needed. This experimental setup closely mirrors real-world conditions and allows for a more thorough evaluation of the tool’s ability to handle dynamic and varied logging requirements, moving beyond the oversimplified assumption that each snippet requires only one log statement. ultimately limiting tool performance in practical applications.

### v3

@@ 4. EXPERIMENTS
  • RQ1: How similar are automatically generated log statements to human-written logging practices?
  • RQ2: How compilable are the log statements generated by automatic logging tools?
  • RQ3: How consistent are runtime logs from auto-logged programs with oracle-generated logs?
Specifically, RQ1 aims to evaluate the performance of the tools in generating log statements that closely resemble the ground truth. RQ2 focuses on assessing the compilability of the predicted log statements, determining whether these predictions can be seamlessly integrated into the code without causing compilation errors. Finally, RQ3 evaluates the runtime logs produced by the generated log statements. We examine the similarity between the predicted logs and the expected logs. We begin by introducing the automatic logging tools selected for evaluation. Then we provide a detailed analysis of these tools by answering the three research questions. This comprehensive evaluation allows us to draw insights into the strengths and limitations of current approaches for automatic logging.
@@ 4.1. Automatic Logging Tools
Automatic logging is a hot topic, leading to the development of many tools for determining specific parts of log statements and end-to-end automatic logging tools in recent years. In this paper, we focus on end-to-end logging tools. We reached out to the authors of popular end-to-end automatic logging tools for assistance in rebuilding these tools. Because SCLogger (Li et al., 2024) is still under construction, we ultimately selected four methods for evaluation: LANCE (Mastropaolo et al., 2022), LEONID (Mastropaolo et al., 2023), FastLog (Xie et al., 2024), UniLog (Xu et al., 2024a). We detailed the method in the following.
LANCE: $LANCE$ (Mastropaolo et al., 2022) is the first model designed to generate and insert complete log statements in code. It takes a method requiring a log statement and outputs a meaningful log message with an appropriate logging level in the correct position. Built on the Text-To-Text Transfer Transformer (T5) model (Raffel et al., 2019), $LANCE$ is trained specifically for injecting proper logging statements.
@@
UniLog: $UniLog$ (Xu et al., 2024a) is the first attempt to adapt Warm-up and In-context-learning strategy to enhance the model’s ability to generate log statements. Due to limitations in assessing the original UniLog, we reproduced it using two backbone models: CodeLlama-7B (Rozière et al., 2023) and DeepSeek-V3 (DeepSeek-AI et al., 2024). We applied the warmup process exclusively to the CodeLlama-7B backbone model while employing the ICL strategy to construct prompts for both models. The data are sourced from LANCE (Mastropaolo et al., 2022) to warm up and generate ICL content. The effectiveness of In-Context Learning often depends on whether the examples are in-distribution or out-of-distribution relative to the evaluation data. Since LANCE’s data distribution differs from our evaluation data, this may affect UniLog’s performance. We will use $\bm{UniLog_{{cl}}}$ to represent the version based on CodeLlama-7B, $\bm{UniLog_{{ds}}}$ to represent the version based on DeepSeek-V3.
@@
FastLog: FastLog (Xie et al., 2024) defines the logging task in two steps: finding the position and generating and inserting a complete log statement into the source code. This approach avoids rewriting the source code, a key limitation of $LANCE$. They utilized PLBART (Ahmad et al., 2021) as the base model to fine-tune two separate models: one for predicting insertion position, the other for generating log statements. With the heuristic rule, log statements only appear after certain special characters, FastLog enhances efficiency while maintaining accuracy in generating log statements.
Table 2. Static evaluation on the complete logging task was conducted. Note that lower ALD means better log level quality, and BLEU-4 & ROUGE-L refer to two kinds of STS metrics. The best results are highlighted.
Method PA LA MA ALD DEA BLEU-4 ROUGE-L
FastLog 58.39 62.15 6.93 0.63 18.00 20.14 29.32
$UniLog_{{ds}}$ $37.11$ $60.66$ $5.23$ 0.61 16.20 11.62 26.37
$UniLog_{{cl}}$ $23.49$ $50.97$ $2.71$ 0.79 16.60 8.79 29.88
LANCE $35.97$ $37.70$ $3.11$ 2.11 15.25 6.7 15.64
$LEONID_{S}$ $11.26$ $17.90$ $1.96$ 3.78 8.51 3.45 6.69
4.2. RQ1: How similar are automatically generated log statements to human-written logging practices?
Table 3. The performance of tools when facing different length input data. Short means data shorter than 512 tokens, Long means data longer than 512 tokens, and $\Delta$ means the difference in logging tools performance when facing the Short data and the Long data.
Method Dataset PA LA MA ALD DEA BLEU-4 ROUGE-L
FastLog Short 62.27 61.79 6.81 0.61 19.65 19.71 29.43
Long 45.04 63.37 7.37 0.68 12.35 21.63 28.9
$\Delta$ $\downarrow 17.23$ $\uparrow 1.58$ $\uparrow 1.58$ $\uparrow 0.56$ $\downarrow 7.30$ $\uparrow 1.92$ $\downarrow 0.47$
$UniLog_{d}{}_{s}$ Short 41.19 59.00 4.11 0.63 16.21 9.93 24.51
Long 23.10 66.38 9.06 0.54 16.19 17.40 32.77
$\Delta$ $\downarrow 18.09$ $\uparrow 7.38$ $\uparrow 4.95$ $\downarrow 0.09$ $\downarrow 0.03$ $\uparrow 7.47$ $\uparrow 8.26$
$UniLog_{c}{}_{l}$ Short 27.78 51.07 2.64 0.77 17.53 8.86 30.24
Long 8.73 50.59 2.94 0.82 13.39 8.44 24.58
$\Delta$ $\downarrow 19.05$ $\downarrow 0.48$ $\uparrow 0.30$ $\uparrow 0.09$ $\downarrow 4.14$ $\downarrow 0.42$ $\downarrow 5.66$
LANCE Short 46.44 48.67 4.01 1.56 19.69 8.65 20.19
Long 0 0 0 0 0 0 0
$\Delta$
$LEONID_{S}$ Short 14.54 23.11 2.53 3.43 10.99 4.45 8.64
Long 0 0 0 0 0 0 0
$\Delta$
We evaluated four automatic logging tools using AL-Bench, leveraging 39,600 static evaluation pairs. The evaluation results are presented in Table 2. All logging tools show average accuracy drops of 37.49%, 23.43%, and 15.80% in predicting log position, level, and message compared to their reported results. Most strikingly, $LEONID_{S}$ collapses from 31.55 → 1.96 in MA ($\downarrow 93.8\%$) and $UniLog_{cl}$ plummets from 76.90 → 23.49 in PA ($\downarrow 69.5\%$). Those drops expose severe overfitting to narrow, synthetic datasets in prior studies. When evaluated on AL-Bench, state-of-the-art tools achieved limited scores of 18.00 (DEA), 20.14 (BLEU-4), and 29.88 (ROUGE-L) in STS. These results highlight persistent limitations in generating semantically coherent natural language descriptions and reliably identifying contextually relevant dynamic content for logging.
To investigate the performance discrepancy relative to prior studies, we partitioned the dataset into two subsets: instances with sub-512 token lengths and those exceeding 512 tokens. This division enables us to determine whether instance length contributes to the performance discrepancies observed among the logging tools, providing further insight into how different data characteristics influence tool effectiveness. After dividing the data, we obtained two groups: long data containing 8,925 instances and short data with 30,675 instances. The evaluation results in Table 3 show a clear difference in tool performance based on the length of the instances. Notably, LANCE and LEONID struggled with instances longer than 512 tokens, failing to generate syntax-correct code and, in some cases, producing incomplete code. This explains why their scores for these cases are reported as zero, highlighting the input length limitations of both tools and their inability to handle longer, more complex instances effectively. $UniLog_{cl}$ and FastLog show a considerable drop in PA and DEA when handling longer data, indicating that they struggle to predict log positions and select dynamic variables in more complex instances. $UniLog_{ds}$ exhibits a significant decline in positional accuracy but only marginal degradation in dynamic variable identification. Notably, it demonstrates stronger performance on metrics such as MA and STS when processing longer input sequences. This observation implies that enhancing a model’s architectural foundation and providing richer contextual data can improve its capacity to generate syntactically and semantically coherent logging statements through greater contextual awareness.
Although $UniLog_{ds}$ is equipped with a more powerful backbone, it still struggles to choose the appropriate position for the log statement. As widely demonstrated that the LLMs are not good at counting numbers (Ahn et al., 2024), making LLM generate the exact line number might not be a wise choice. Furthermore, by simply analyzing the control flow graph of the code, we might be able to exclude positions where logging is not feasible to leave fewer choices.
@@ 4.2. RQ1: How similar are automatically generated log statements to human-written logging practices?
To evaluate tool stability, we analyzed performance consistency across individual projects. As shown in Figure 5, tool performance exhibits significant variability across projects. For instance, in Flink, the state-of-the-art tool achieves an MA score of 13.68, contrasting sharply with the overall project average of 6.93. Similarly, in Keycloak, the state-of-the-art DEA reaches 43.84 compared to the aggregate average of 16.60. These stark discrepancies underscore inconsistencies in the reliability and generalizability of logging tools when applied to projects with distinct requirements and contexts.
Answer to RQ1. All logging tools show average accuracy drops of 37.49%, 23.43%, and 15.80% in predicting log position, level, and message compared to their reported results. All logging tools exhibit persistent limitations in generating semantically coherent log descriptions and reliably identifying contextually relevant dynamic variables. Moreover, they fail to achieve consistent performance reliability across diverse logging requirements and project-specific contexts.
@@
Failed Reason # Failed
@@
4.3. RQ2: How compilable are the log statements generated by automatic logging tools?
To evaluate the tools’ ability to generate compilable log statements, we replace existing log statements with predicted ones and recompile the project to check for successful compilation. As shown in Table 5, the best-performing tool, FastLog, achieves a 79.9% Compilation Success Rate, followed by $UniLog_{c}l$ at 70.3%, $UniLog_{d}s$ at 60.2%, LANCE at 49.4%, $LEONID_{M}$ and $LEONID_{S}$ at 25.0% and 16.4%, respectively. The results reveal a key limitation of LANCE and LEONID that they regenerate the entire code snippet, which increases the risk of unintended code changes and can potentially lead to compilation errors. In contrast, FastLog and UniLog focus solely on generating the new log statement, minimizing the risk of errors by limiting codebase modifications. Although FastLog is one of the best tools available, it still leads to significant instances (20.1%) of compile failure.
To understand the reason for the compilation failures reason, we conducted a manual review of the compilation failures to identify the specific causes. We randomly selected 100 failed instances from the best tool, FastLog, and the first two authors cross-checked the causes. The analysis results are presented in Table 5. The most common failure, occurring in 56 instances, is due to using the wrong logging name (i.e., Logger, LOG), indicating that incorrect or non-existent log functions are being invoked. Using undefined methods accounts for 21 failures, followed by using undefined variables with 15 failures. Less frequent issues include incompatible types (3), and unreachable statements (1). Others mean generating the wrong syntax code, which we will not analyze.
The majority of failures (totaling 92 instances) involve undefined references to methods (21 instances), variables (15 instances), or logging names (56 instances). These undefined reference failures are primarily due to the limited context provided by existing function-level logging tools, lacking critical details on valid variables, methods, libraries, and packages relevant to the target function. Current tools are designed for function-level input, highlighting the need for logging tools to integrate better context awareness and validation checks to ensure compatibility with the existing codebase. The less frequent errors are also noticeable. For instance, the issue of unreachable statements points to a weakness in that tools lack an understanding of the code’s control flow, resulting in log statements placed in non-executable paths. Collectively, these errors highlight the need for automated logging tools to adopt enhanced context-aware strategies, improving the accuracy and contextual relevance of generated logs.
@@ 4.3. RQ2: How compilable are the log statements generated by automatic logging tools?
4.4. RQ3: How consistent are runtime logs from auto-logged programs with oracle-generated logs?
To answer RQ3, we compare the semantic similarity of logs generated by predicted log statements and by the oracle. It is important to highlight that, aside from $LEONID_{M}$, all other tools operate under the strong assumption that the given code snippet requires exactly one log statement. To highlight the inappropriateness of this assumption, we only allow the tools one chance to predict the log statement, even in cases where multiple log statements are needed. Our experimental design emulates real-world logging scenarios by abandoning the assumption that single log statements universally suffice, thereby exposing a critical limitation in existing tools’ capacity to determine contextually appropriate logging density for a given code context.